one x-intercept for a parabola is at the point (3,0).Use the quadratic formula to find the other x-intercept for the parabola defined by the equation: y= x^2-2x-3.
Answer by CharlesG2(834)

one x-intercept for a parabola is at the point (3,0).Use the quadratic formula to find the other x-intercept for the parabola defined by the equation: y= x^2-2x-3.

y = x^2 - 2x - 3
y = (x + 1)(x - 3) by FOIL --> x^2 - 3x + x - 3 (same as above)
x-intercepts: (3,0) and (-1,0)